# Nearly-Linear Time Seeded Extractors with Short Seeds

### v1

@@ 5.2.1 The (extremely) low-error case
It remains to analyze the time complexity and the overall error of the recursive procedure above. Evaluating $\mathsf{Ext}_{i(k)}$ requires at most eight evaluations of the condenser from Theorem 3.11, four evaluations of the fast hash-based extractor from Lemma 2.25, four evaluations of $\mathsf{Ext}_{i(k^{\prime\prime})}$ for some $i(k^{\prime\prime})<i(k)$, and simple operations that can be done in time $\widetilde{O}(n)$. This means that the overall time complexity is $4^{i(k)}\cdot\widetilde{O}(n)=\widetilde{O}(n)$ after a one-time preprocessing step independent of the source and seed, since $4^{i(k)}=\operatorname{poly}(\log n)$ by Equation 2. This preprocessing step corresponds to finding primitive elements for $O(\log\log n)$ fields $\mathds{F}_{q}$ with orders $q\leq\operatorname{poly}(n/\varepsilon_{0})=\operatorname{poly}(n/\varepsilon)$ powers of $2$. Furthermore, $\varepsilon_{i(k)}=O(\varepsilon_{0}+\varepsilon_{i(k/3)})$ for all $k$, and so $\varepsilon_{i(k)}=2^{O(i(k))}\varepsilon_{0}=\operatorname{poly}(\log n)\cdot\varepsilon_{0}\leq\varepsilon$ provided that $\varepsilon_{0}\leq\varepsilon/\log^{C}n$ for a large enough constant $C>0$. ∎
@@ 5.2.2 The (relatively) high-error case
In this section, we consider the higher error case where $\varepsilon\geq Cn^{3}\cdot 2^{-k/\log k}$. We instantiate the recursive approach of Srinivasan and Zuckerman [SZ99, Section 5.5] appropriately with our building blocks and analyze its complexity.
@@
    If $\mathsf{Ext}_{0}$ is computable in time $\widetilde{O}(n_{0})$ and $k\geq C\log^{2}(n/\varepsilon)$, then $\mathsf{Ext}$ is computable in time $\widetilde{O}(n)$;
@@
    If $\mathsf{Ext}_{0}$ is computable in time $\widetilde{O}(n_{0})$ after a preprocessing step corresponding to finding primitive elements of $j$ fields $\mathds{F}_{q}$ of orders $q\leq\operatorname{poly}(n/\varepsilon_{0})$, then $\mathsf{Ext}$ is computable in time $\widetilde{O}(n)$ after a preprocessing step corresponding to finding primitive elements of $j+1$ fields $\mathds{F}_{q}$ of orders $q\leq\operatorname{poly}(n/\varepsilon_{0})$.
@@
We begin by setting up relevant notation:
@@
  • Set $t=1+\frac{\log\mathopen{}\mathclose{{}\left(u/\log u}\right)}{\log(1/0.9)}$. This will be the number of blocks of $Z$. We have $m_{t}=0.9^{t-1}m_{1}=\log u\cdot\log(\ell_{0}/\varepsilon_{0})$. Furthermore, since $\ell_{1}=C_{\mathsf{blocks}}\cdot m_{1}$, we also have that $\ell_{i}=C_{\mathsf{blocks}}\cdot m_{i}$ for all $i\geq 1$.
@@
    Using a fresh seed $Y_{\mathsf{Cond}}$ of length $C_{\mathsf{Cond}}\log(n/\varepsilon_{0})$, apply a strong $(k,k^{\prime},\varepsilon_{0}^{2})$-condenser $\mathsf{Cond}$ to $X$. If $k\geq C\log^{2}(n/\varepsilon_{0})$ for an apropriately large constant $C>0$, then we instantiate $\mathsf{Cond}$ with the lossless KT $(k,k^{\prime}=k,\varepsilon_{\mathsf{Cond}})$-condenser (Theorem 3.10). Otherwise, we instantiate $\mathsf{Cond}$ with the lossy RS $(k,k^{\prime}\geq 0.95k,\varepsilon_{\mathsf{Cond}})$-condenser (Theorem 3.11) instantiated with $\alpha=0.05$. By the second part of either Theorem 3.10 or Theorem 3.11, we get that with probability at least $1-\varepsilon_{0}$ over the choice of $Y_{\mathsf{Cond}}=y$ it holds that $X^{\prime}=\mathsf{Cond}(X,y)$ is $\varepsilon_{0}$-close to an $(n^{\prime},k^{\prime})$-source with $k^{\prime}\geq 0.95n^{\prime}$. From here onwards we work under such a good fixing $Y_{\mathsf{Cond}}=y$, and will account for the $\varepsilon_{0}$ error term in the final extractor error later on.
@@
    We use $X^{\prime}$ and Lemma 2.16 to generate a block source $Z$ with geometrically decreasing block lengths $\ell_{0},\ell_{1},\dots,\ell_{t}$ defined above.
    For each $i=0,1,\dots,t$, let $\mathsf{Samp}_{i}\colon\{0,1\}^{r_{i}}\to[n^{\prime}]^{\ell_{i}}$ be the $(\theta=1/100,\gamma=\varepsilon_{0})$-averaging sampler from Lemma 2.21 with input length $r_{i}=C_{\mathsf{Samp}}\log(n^{\prime}/\varepsilon_{0})$ for some constant $C_{\mathsf{Samp}}>0$. We choose the constant $C_{\mathsf{blocks}}$ above to be large enough so that $n^{\prime}\geq\ell_{i}\geq\ell_{t}\geq C^{\prime}_{\mathsf{Samp}}\log(1/\varepsilon_{0})/\theta^{2}$ for all $i\in[t]$, where $C^{\prime}_{\mathsf{Samp}}$ is the constant $C$ from Lemma 2.21. To see that $\ell_{i}\leq n^{\prime}$ for $i=0,\dots,t$ (and so indeed Lemma 2.21 can be applied to obtain $\ell_{i}$ samples), note that
    $\sum_{i=0}^{t}\ell_{i}\leq\sum_{i=0}^{\infty}\ell_{i}=10\ell_{1}+\ell_{0}\leq k/9<n^{\prime}.$ (6)
@@
    where the first inequality holds since $u\log(\ell_{0}/\varepsilon_{0})\leq k_{0}$ is an hypothesis in the lemma statement. We also assume that $\varepsilon_{0}\geq 2^{-ck_{0}}$ for a constant $c>0$ small enough so that
    $\ell_{0}=\frac{k}{100C_{\mathsf{blocks}}}\geq C^{\prime}_{\mathsf{Samp}}\cdot ck_{0}/\theta^{2}\geq C^{\prime}_{\mathsf{Samp}}\log(1/\varepsilon_{0})/\theta^{2},$
    where we recall that $k_{0}=0.7\ell_{0}$, meaning that the conditions of Lemma 2.21 are satisfied for all $i=0,\dots,t$.
@@
    For each $i=0,1,\dots,t$, let $Y_{i}$ be a fresh seed of length $r_{i}$. We set the $i$-th block as $Z_{i}=X^{\prime}_{\mathsf{Samp}(Y_{i})}$. By Lemma 2.16 instantiated with $X^{\prime}$ and $\mathsf{Samp}_{0}$, we conclude that
    $Y_{0}\circ Z_{0}\approx_{\varepsilon_{0}+2^{-c_{\mathsf{Samp}}k^{\prime}}}Y_{0}\circ Z^{\prime}_{0},$
@@
    with $c_{\mathsf{Samp}}>0$ an absolute constant guaranteed by Lemma 2.16, where $(Z^{\prime}_{0}|Y_{0}=y)$ is an $(\ell_{0},0.9\ell_{0})$-source for every $y$. We now argue how this guarantee extends to more blocks. Consider an arbitrary $i$ and fixings $Y_{0}=y_{0},\dots,Y_{i-1}=y_{i-1}$. Then, Lemma 2.6 with $\delta=2^{-c_{\mathsf{Samp}}k}$ and $\ell=k/9$ (from the upper bound in Equation 6) implies that
    $\mathbf{H}_{\infty}(X|(Z^{\prime}_{0}|Y_{0}=y_{0})=z_{0},\dots,(Z^{\prime}_{i-1}|Y_{i-1}=y_{i-1})=z_{i-1})\geq 0.8n^{\prime}$
@@
    except with probability at most $2^{-c_{\mathsf{Samp}}k}$ over the choice of $z_{0},\dots,z_{i-1}$, which we can absorb into the statistical distance, since $k^{\prime}\geq 0.95n^{\prime}\geq 0.95k$. Consequently, from Lemma 2.16 we get that
@@
    $Y_{0},Z^{\prime}_{0},\dots,Y_{i-1},Z^{\prime}_{i-1},Y_{i},Z_{i}=X_{\mathsf{Samp}(Y_{i})}\approx_{\varepsilon_{0}+2\cdot 2^{-c_{\mathsf{Samp}}k}}Y_{0},Z^{\prime}_{0},\dots,Y_{i-1},Z^{\prime}_{i-1},Y_{i},Z^{\prime}_{i},$ (7)
@@
    where $(Z^{\prime}_{i}|Y_{0}=y_{0},Z^{\prime}_{0}=z_{0},\dots,Y_{i-1}=y_{i-1},Z^{\prime}_{i-1}=z_{i-1},Y_{i}=y_{i})$ is an $(\ell_{i},0.7\ell_{i})$-source for any choice of $y_{0},z_{0},\dots,y_{i-1},z_{i-1},y_{i}$. Combining Equation 7 with the triangle inequality over all $0\leq i\leq t$, we conclude that $Z=Z_{0}\circ Z_{1}\circ\cdots\circ Z_{t}$ is $\varepsilon_{\mathsf{block}}$-close to an exact $(\ell_{0},\dots,\ell_{t},0.7)$-block-source $Z^{\prime}$, where $\varepsilon_{\mathsf{block}}=(t+1)(\varepsilon_{0}+2\cdot 2^{-c_{\mathsf{Samp}}k})$.
@@
    We apply block source extraction (Lemma 2.23) to $Z=Z_{0}\circ Z_{1}\circ\cdots\circ Z_{t}$. More precisely, let $\mathsf{BExt}\colon\{0,1\}^{\ell_{0}}\times\cdots\times\{0,1\}^{\ell_{t}}\times\{0,1\}^{d_{t}}\to\{0,1\}^{m_{0}}$ be the strong $(k_{0},k_{1},\dots,k_{t},(t+1)\varepsilon_{0})$-block-source extractor with $k_{i}=0.7\ell_{i}$ obtained via Lemma 2.23 as follows. We instantiate $\mathsf{Ext}_{0}$ with the strong extractor promised by the lemma statement with seed length $d_{0}\leq u\cdot\log(\ell_{0}/\varepsilon_{0})=m_{1}$. For $i\in[t]$, we instantiate $\mathsf{Ext}_{i}\colon\{0,1\}^{\ell_{i}}\times\{0,1\}^{d_{i}}\to\{0,1\}^{m_{i}}$ as the strong $(k_{i}=0.7\ell_{i},\varepsilon_{0})$-seeded extractor from Lemma 2.14 with seed length $d_{i}=2m_{i}+4\log(\ell_{i}/\varepsilon_{0})+8$. We choose the constant $C_{\mathsf{blocks}}$ to be large enough so that
@@
    as required by Lemma 2.14. This is possible since by choosing $C_{\mathsf{blocks}}$ large enough we have
    $\ell_{i}\geq\ell_{t}=C_{\mathsf{blocks}}\cdot m_{t}=C_{\mathsf{blocks}}\log u\cdot\log(\ell_{0}/\varepsilon_{0})\geq 100\log(4/\varepsilon_{0})$
    for all $i\in[t]$, and so $20.7\ell_{i}-16\log(4/\varepsilon_{0})\geq\ell_{i}/$ 2 for all $i\in[t]$. Furthermore, for any $i\geq 2$ the output length $m_{i}$ of $\mathsf{Ext}_{i}$ satisfies
@@
    where we recall that $m_{i}=m_{i-1}/0.9$ for $i\geq 2$. Finally, the output length of $\mathsf{Ext}_{1}$ satisfies $d_{1}+m_{1}\geq m_{1}\geq d_{0}$, where we recall that $d_{0}$ is the seed length of $\mathsf{Ext}_{0}$.
@@
    since $\ell_{0}\leq k\leq n$. By Lemma 2.23, we get that
    $Y_{\mathsf{BExt}}\circ\mathsf{BExt}(Z,Y_{\mathsf{BExt}})\approx_{\varepsilon_{\mathsf{block}}}Y_{\mathsf{BExt}}\circ\mathsf{BExt}(Z^{\prime},Y_{\mathsf{BExt}})\approx_{(t+1)\varepsilon_{0}}U_{d_{t}+m_{0}}.$

### v2

@@ 4.2.1 The (extremely) low-error case
It remains to analyze the time complexity and the overall error of the recursive procedure above. Evaluating $\mathsf{Ext}_{i(k)}$ requires at most eight evaluations of the condenser from theorem˜3.17, four evaluations of the fast hash-based extractor from lemma˜2.24, four evaluations of $\mathsf{Ext}_{i(k^{\prime\prime})}$ for some $i(k^{\prime\prime})<i(k)$, and simple operations that can be done in time $\widetilde{O}(n)$. This means that the overall time complexity is $4^{i(k)}\cdot\widetilde{O}(n)=\widetilde{O}(n)$ after a one-time preprocessing step independent of the source and seed, since $4^{i(k)}=\operatorname{poly}(\log n)$ by eq.˜10. This preprocessing step corresponds to finding primitive elements for $O(\log\log n)$ fields $\mathds{F}_{q}$ with prime orders $q\leq\operatorname{poly}(n/\varepsilon_{0})=\operatorname{poly}(n/\varepsilon)$. Furthermore, $\varepsilon_{i(k)}=O(\varepsilon_{0}+\varepsilon_{i(k/3)})$ for all $k$, and so $\varepsilon_{i(k)}=2^{O(i(k))}\varepsilon_{0}=\operatorname{poly}(\log n)\cdot\varepsilon_{0}\leq\varepsilon$ provided that $\varepsilon_{0}\leq\varepsilon/\log^{C}n$ for a large enough constant $C>0$. ∎
@@ 4.2.2 The (relatively) high-error case
In this section, we consider the higher error case where $\varepsilon\geq Cn^{3}\cdot 2^{-k/\log k}$. We instantiate the recursive approach of Srinivasan and Zuckerman [SZ99, Section 5.5] appropriately with the fast condensers from section˜3.3, the sampler from lemma˜2.20, and the fast hash-based seeded extractors from lemma˜2.13, and analyze its complexity.
The next lemma shows how we can recursively decrease the seed length of an extractor. We complete the construction by instantiating the base extractor in this recursion appropriately, and then increasing its output length.
@@
    If $\mathsf{Ext}_{0}$ is computable in time $T(n_{0})$ and $k\geq C\log^{2}(n/\varepsilon_{0}(ck))$, then $\mathsf{Ext}$ is computable in time $T(n)+\widetilde{O}(n+\sqrt{n}\cdot\log(1/\varepsilon_{0}(ck))^{5})$;
@@
    If $\mathsf{Ext}_{0}$ is computable in time $T(n_{0})$ after a preprocessing step, then $\mathsf{Ext}$ is computable in time $T(n)+\widetilde{O}(n)$ after a preprocessing step. [^27] We discuss the precise preprocessing step in more detail in remark 4.11.
@@
We begin by discussing the high-level approach in this proof. On input an arbitrary $(n,k)$-source $X$, $\mathsf{Ext}$ proceeds as follows. First, it applies a fast strong condenser to $X$ to obtain a new source $X^{\prime}$ with high min-entropy rate. If $k\geq C\log^{2}(n/\varepsilon_{0})$ then we can apply the KT condenser, which does not require preprocessing unless $\varepsilon_{0}$ is very small. Otherwise, we apply the RS condenser. Then, we use $X^{\prime}$ to generate a block source $Z=(Z_{0},\dots,Z_{t})$ with geometrically decreasing block lengths. The way we achieve this depends on the regime we are in. If we are in a regime where we must anyway resort to the RS condenser, then we generate each block by applying an appropriately instantiated RS condenser with a fresh seed to $X^{\prime}$. Otherwise, if we are in a regime where we can use the KT condenser, then we use the expander random walks averaging sampler from lemma˜2.20, which runs in time $\widetilde{O}(n)$ in this regime. [^28] The sole reason for this case analysis is that by using the sampler instead of the RS condenser in the “KT regime” we can avoid a preprocessing step unless $\varepsilon$ is tiny. Finally, we apply block source extraction to $Z$. More concretely, we begin by applying the fast hash-based extractor from lemma˜2.13 to the shorter blocks at the end of $Z$, up until the second block $Z_{1}$ of $Z$. This generates a sufficiently large (but still short) seed that we can use to extract from the first block $Z_{0}$ using the base extractor $\mathsf{Ext}_{0}$.
We now formally analyze the approach above. We begin by setting up relevant parameters:
@@
  • Set $t=1+\mathopen{}\mathclose{{\left\lceil\frac{\log\mathopen{}\mathclose{{\left(u/\log u}}\right)}{\log(1/0.9)}}}\right\rceil$. This will be the number of blocks of $Z$. We have $m_{t}=0.9^{t-1}m_{1}\in[0.9\log u\cdot\log(\ell_{0}/\varepsilon_{0}),\log u\cdot\log(\ell_{0}/\varepsilon_{0})]$. Furthermore, since $\ell_{1}=C_{\mathsf{blocks}}\cdot m_{1}$, we also have that $\ell_{i}=C_{\mathsf{blocks}}\cdot m_{i}$ for all $i\geq 1$.
@@
    Using a fresh seed $Y_{\mathsf{Cond}}$ of length $C_{\mathsf{Cond}}\log(n/\varepsilon_{0})$, apply a strong $(k,k^{\prime},\varepsilon_{0}^{2})$-condenser $\mathsf{Cond}$ to $X$. If $k\geq C\log^{2}(n/\varepsilon_{0})$ for an apropriately large constant $C>0$, then we instantiate $\mathsf{Cond}$ with the KT strong $(k,k^{\prime}=k,\varepsilon_{0}^{2})$-condenser (theorem˜3.14) instantiated with $\alpha=0.05$. Otherwise, we instantiate $\mathsf{Cond}$ with the lossy RS $(k,k^{\prime}\geq 0.975k,\varepsilon_{0}^{2})$-condenser (theorem˜3.17) instantiated with $\alpha=0.025$. By the second part of either theorem˜3.14 or theorem˜3.17, we get that with probability at least $1-\varepsilon_{0}$ over the choice of $Y_{\mathsf{Cond}}=y$ it holds that $X^{\prime}=\mathsf{Cond}(X,y)$ is $\varepsilon_{0}$-close to an $(n^{\prime},k^{\prime})$-source with $k^{\prime}\geq 0.95n^{\prime}$.
    From here onwards we work under such a good fixing $Y_{\mathsf{Cond}}=y$ and also assume that $X^{\prime}$ is an $(n^{\prime},k^{\prime})$-source. We account for the resulting $2\varepsilon_{0}$ error term in the final extractor error later.
@@
    We use $X^{\prime}$ to generate a block source $Z=(Z_{0},Z_{1},\dots,Z_{t})$ with geometrically decreasing block lengths. Our procedure depends on the regime of parameters we are in:
    1. (a)
      If $k\geq C\log^{2}(n/\varepsilon_{0})$ for an appropriately large constant $C>0$, then for each $i=0,1,\dots,t$ let $\mathsf{Samp}_{i}\colon\{0,1\}^{r_{i}}\to[n^{\prime}]^{\ell_{i}}$ be the $(\gamma=\varepsilon_{0},\theta=1/100)$-averaging sampler from lemma˜2.20 with input length $r_{i}=\gamma_{\mathsf{Samp}}\log(n^{\prime}/\varepsilon_{0})$ for some constant $\gamma_{\mathsf{Samp}}>0$. We choose the constant $C_{\mathsf{blocks}}$ above to be large enough so that $n^{\prime}\geq\ell_{i}\geq\ell_{t}\geq C_{\mathsf{Samp}}\log(1/\varepsilon_{0})/\theta^{2}$ for all $i\in[t]$, where $C_{\mathsf{Samp}}$ is the constant $C$ from lemma˜2.20. To see that $\ell_{i}\leq n^{\prime}$ for $i=0,1,\dots,t$ (and so indeed lemma˜2.20 can be applied to obtain $\ell_{i}$ samples), note that
      $\ell_{0}+\sum_{i=1}^{t}\ell_{i}\leq\ell_{0}+\sum_{i=1}^{\infty}\ell_{i}=\ell_{0}+10\ell_{1}\leq k/9<n^{\prime}.$ (14)
@@
      where the first inequality holds since $u\log(\ell_{0}/\varepsilon_{0})\leq k_{0}$ is an hypothesis in the lemma statement.
@@
      By lemma˜2.15 instantiated with $X^{\prime}$ and $\mathsf{Samp}_{0}$, we conclude that
      $(Y_{0},Z_{0})\approx_{\varepsilon_{0}+2^{-\beta_{\mathsf{Samp}}k}}(Y_{0},Z^{\prime}_{0}),$ (15)
@@
      with $\beta_{\mathsf{Samp}}>0$ the constant guaranteed by lemma˜2.15, where $(Z^{\prime}_{0}|Y_{0}=y_{0})$ is an $(\ell_{0},0.9\ell_{0})$-source for every $y_{0}$. We now argue how this guarantee extends to more blocks.
      For each $Z_{j}$, define $Z_{j,\vec{y}}=(Z_{j}|(Y_{0},\dots,Y_{i-1})=\vec{y})$. Consider any fixing $(Y_{0},\dots,Y_{i-1})=\vec{y}$. Then, lemma˜2.6 with $\delta=2^{-\beta_{\mathsf{Samp}}k}$, where $\beta_{\mathsf{Samp}}>0$ is taken to be a small enough constant, and $\ell=k/9$ (from the upper bound in eq.˜14) implies that
      $\mathbf{H}_{\infty}(X^{\prime}|Z_{0,\vec{y}}=z_{0},\dots,Z_{i-1,\vec{y}}=z_{i-1})$ $\geq k^{\prime}-\ell-\beta_{\mathsf{Samp}}k$
      $\geq 0.95n^{\prime}-k/9-\beta_{\mathsf{Samp}}k$
      $\geq 0.95n^{\prime}-n^{\prime}/9-\beta_{\mathsf{Samp}}n^{\prime}$
      $\geq 0.8n^{\prime}$ (16)
@@
      except with probability at most $2^{-\beta_{\mathsf{Samp}}k}$ over the choice $(z_{0},\dots,z_{i-1})\sim(Z_{0,\vec{y}},\dots,Z_{i-1,\vec{y}})$. Call a fixing $\vec{v}=(y_{0},z_{0},\dots,y_{i-1},z_{i-1})$ for which item˜2a holds good. Define $X^{\prime}_{\vec{v}}=(X^{\prime}|(Y_{0},Z_{0},\dots,Y_{i-1},Z_{i-1})=\vec{v})$. Then, by item˜2a and lemma˜2.15 we know that for all good $\vec{v}$-s we have
@@
      $\mathopen{}\mathclose{{\left(Y_{i},Z_{i,\vec{v}}=(X^{\prime}_{\vec{v}})_{\mathsf{Samp}_{i}(Y_{i})}}}\right)\approx_{\varepsilon_{0}+2^{-\beta_{\mathsf{Samp}}k}}\mathopen{}\mathclose{{\left(Y_{i},Z^{\prime}_{i,\vec{v}}}}\right),$ (17)
      with $(Z^{\prime}_{i,\vec{v}}|Y_{i}=y_{i})$ an $(\ell_{i},0.7\ell_{i})$-source for all $y_{i}$.
      Analogously to [NZ96, proof of Lemma 17] and the proof of lemma˜4.4, we can use eq.˜17 and the fact that $\vec{v}\sim(Y_{0},Z_{0,\vec{y}},\dots,Y_{i-1},Z_{i-1,\vec{y}})$ is good with probability at least $1-2^{-\beta_{\mathsf{Samp}}k}$ to show by induction on the number of blocks that
      $(Y_{0},\dots,Y_{t},Z)\approx_{\varepsilon_{\mathsf{block}}}(Y_{0},\dots,Y_{t},Z^{\prime\prime}),$ (18)
@@
      where for every $(y_{0},\dots,y_{t})$, $(Z^{\prime\prime}|Y_{0}=y_{0},\dots,Y_{t}=y_{t})$ is an exact $((\ell_{0},\dots,\ell_{t}),0.7)$-block-source, and $\varepsilon_{\mathsf{block}}=(t+1)(\varepsilon_{0}+2\cdot 2^{-\beta_{\mathsf{Samp}}k})$.
    2. (b)
      If $k<C\log^{2}(n/\varepsilon_{0})$, then for each $i=0,1,\dots,t$ let $\mathsf{Cond}_{i}\colon\{0,1\}^{n^{\prime}}\to\{0,1\}^{m_{i}}$ be the strong RS $(k_{i}=0.9\ell_{i},k^{\prime}_{i},\varepsilon_{0}^{2})$-condenser from theorem˜3.17 instantiated with $\alpha=0.01$. Note that $k_{i}\leq m_{i}\leq(1+\alpha)k_{i}\leq\ell_{i}$ and $k^{\prime}_{i}\geq(1-\alpha)m_{i}\geq 0.99k_{i}$. Using a fresh seed $Y_{i}$ of length at most $C^{\prime}_{\alpha}\log(n^{\prime}/\varepsilon_{0}^{2})\leq 2C^{\prime}_{\alpha}\log(n^{\prime}/\varepsilon_{0})$ with $C^{\prime}_{\alpha}$ the constant guaranteed by theorem˜3.17 for $\alpha=0.01$, we compute $W_{i}=\mathsf{Cond}_{i}(X^{\prime},Y_{i})$ and obtain $Z_{i}$ by padding $W_{i}$ to get length exactly $\ell_{i}$, i.e., $Z_{i}=(W_{i},0^{\ell_{i}-m_{i}})$. This is valid since, as already pointed out, $m_{i}\leq\ell_{i}$. Later we argue that, despite this padding, $Z_{i}$ will be statistically close to a source $Z^{\prime}_{i}$ with sufficiently large min-entropy rate.
      The argument showing that $Z_{0},Z_{1},\dots,Z_{t}$ is close to an exact $((\ell_{0},\ell_{1},\dots,\ell_{t}),0.7)$-block-source conditioned on the seeds $Y_{1},\dots,Y_{t}$ is very similar to that of the previous case. Nevertheless, we do need to check that the choices of $\ell_{0},\dots,\ell_{t}$ allow the desired applications of theorem˜3.17.
      First, to apply theorem˜3.17 we need that $k_{i}\geq C_{\alpha}\log(n^{\prime}/\varepsilon_{0}^{2})$ for all $i$, with $C_{\alpha}>0$ the constant from theorem˜3.17 for $\alpha=0.01$. Since $k_{i}=0.9\ell_{i}\geq 0.9\ell_{t}=k_{t}$ for all $i$, it suffices to show that $k_{t}\geq C_{\alpha}\log(n^{\prime}/\varepsilon_{0}^{2})$. Since $n^{\prime}\leq 2k$ and
      $k_{t}=0.9\ell_{t}=0.9^{t}\ell_{1}\geq 0.9^{2}C_{\mathsf{blocks}}\log(\ell_{0}/\varepsilon_{0})=0.9^{2}C_{\mathsf{blocks}}\log\mathopen{}\mathclose{{\left(\frac{k}{100C_{\mathsf{blocks}}\varepsilon_{0}}}}\right),$
      it is enough to guarantee that
      $0.9^{2}C_{\mathsf{blocks}}\log\mathopen{}\mathclose{{\left(\frac{k}{100C_{\mathsf{blocks}}\varepsilon_{0}}}}\right)\geq 2C_{\alpha}\log(2k/\varepsilon_{0}).$
      If we take $C_{\mathsf{blocks}}$ to be large enough so that $C_{\mathsf{blocks}}\geq 5C_{\alpha}$, the desired inequality holds provided that, say, $k\geq 2(100C_{\mathsf{blocks}})^{2}$, which is a constant lower bound on $k$.
      Analogously to item˜2a in the previous case, for any $i=0,1,\dots,t$ and an arbitrary fixing $\vec{y}=(y_{0},\dots,y_{i-1})$ we have that
      $\mathbf{H}_{\infty}(X^{\prime}|Z_{0,\vec{y}}=z_{0},\dots,Z_{i-1,\vec{y}}=z_{i-1})\geq 0.8n^{\prime}$
      except with probability at most $2^{-\beta k}$ over the choice of $z_{0},\dots,z_{i-1}$, for a sufficiently small constant $\beta>0$. Therefore, under such a good fixing $\vec{v}=(y_{0},z_{0},\dots,y_{i-1},z_{i-1})$, and defining $W_{i,\vec{v}}$ to be $W_{i}$ conditioned on $(Y_{0},Z_{0},\dots,Y_{i-1},Z_{i-1})=\vec{v}$, theorem˜3.17 guarantees that $(Y_{i},W_{i,\vec{v}})\approx_{\varepsilon_{0}}(Y_{i},W^{\prime}_{i,\vec{v}})$ with $(W^{\prime}_{i,\vec{v}}|Y_{i}=y_{i})$ an $(m_{i},k^{\prime}_{i})$-source for all $y_{i}$ provided that $0.8n^{\prime}\geq k_{i}=0.9\ell_{i}$. This holds, since
      $0.8n^{\prime}\geq 0.8k\geq\frac{k}{100C_{\mathsf{blocks}}}=\ell_{0}\geq 0.9\ell_{i}=k_{i}$
      for any $i=0,1,\dots,t$. By padding $W_{i,\vec{v}}$ and $W^{\prime}_{i,\vec{v}}$ with $0^{\ell_{i}-m_{i}}$, we get that $(Y_{i},Z_{i,\vec{v}})\approx_{\varepsilon_{0}}(Y_{i},Z^{\prime}_{i,\vec{v}})$ with $(Z^{\prime}_{i,\vec{v}}|Y_{i}=y_{i})$ an $(\ell_{i},k^{\prime}_{i})$-source for all $y_{i}$. Furthermore, the min-entropy of $Z^{\prime}_{i,\vec{v}}$ satisfies
      $k^{\prime}_{i}\geq(1-\alpha)k_{i}=0.99k_{i}=0.99\cdot 0.9\ell_{i}\geq 0.7\ell_{i},$
      and so $(Z^{\prime}_{i,\vec{v}}|Y_{i}=y_{i})$ is an $(\ell_{i},0.7\ell_{i})$-source for all $y_{i}$. As in the previous case, this can be used to conclude by induction that
      $(Y_{0},\dots,Y_{t},Z)\approx_{\varepsilon^{\prime}_{\mathsf{block}}}(Y_{0},\dots,Y_{t},Z^{\prime\prime}),$
      where for every fixing $(y_{0},\dots,y_{t})$ we have that $(Z^{\prime\prime}|Y_{0}=y_{0},\dots,Y_{t}=y_{t})$ is an exact $((\ell_{0},\dots,\ell_{t}),0.7)$-block-source, and $\varepsilon^{\prime}_{\mathsf{block}}=(t+1)(\varepsilon_{0}+2^{-\beta k})$. We choose $\beta=\beta_{\mathsf{Samp}}$ as in item˜2a, and so $\varepsilon^{\prime}_{\mathsf{block}}\leq\varepsilon_{\mathsf{block}}$.
@@
    We apply block source extraction (lemma˜2.22) to $Z=(Z_{0},Z_{1},\dots,Z_{t})$. More precisely, let $\mathsf{BExt}\colon\{0,1\}^{\ell_{0}}\times\cdots\times\{0,1\}^{\ell_{t}}\times\{0,1\}^{d_{t}}\to\{0,1\}^{m_{0}}$ be the strong $(k_{0},k_{1},\dots,k_{t},(t+1)\varepsilon_{0})$-block-source extractor with $k_{i}=0.7\ell_{i}$ obtained via lemma˜2.22 as follows. We instantiate $\mathsf{Ext}_{0}$ with the strong extractor promised by the lemma statement with seed length $d_{0}\leq u\cdot\log(\ell_{0}/\varepsilon_{0})=m_{1}$. For $i\in[t]$, we instantiate $\mathsf{Ext}_{i}\colon\{0,1\}^{\ell_{i}}\times\{0,1\}^{d_{i}}\to\{0,1\}^{m_{i}}$ as the strong $(k_{i}=0.7\ell_{i},\varepsilon_{0})$-seeded extractor from lemma˜2.13 with seed length $d_{i}=2m_{i}+4\log(\ell_{i}/\varepsilon_{0})+8$. We choose the constant $C_{\mathsf{blocks}}$ to be large enough so that
@@
    as required by lemma˜2.13. This is possible since by choosing $C_{\mathsf{blocks}}$ large enough we have
    $\ell_{i}\geq\ell_{t}=C_{\mathsf{blocks}}\cdot m_{t}\geq 0.9\cdot C_{\mathsf{blocks}}\log u\cdot\log(\ell_{0}/\varepsilon_{0})\geq 100\log(4/\varepsilon_{0})$
    for all $i\in[t]$, and so $0.7\ell_{i}-16\log(4/\varepsilon_{0})\geq\ell_{i}/2$ for all $i\in[t]$. Furthermore, for any $i\geq 2$ the output length $m_{i}$ of $\mathsf{Ext}_{i}$ satisfies
@@
    where we recall that $m_{i}=0.9m_{i-1}$ for $i\geq 2$. Finally, the output length of $\mathsf{Ext}_{1}$ satisfies $d_{1}+m_{1}\geq m_{1}\geq d_{0}$, where we recall that $d_{0}$ is the seed length of $\mathsf{Ext}_{0}$.
@@
    since $\ell_{0}\leq k\leq n$. By lemma˜2.22, we get that
    $\mathopen{}\mathclose{{\left(Y_{0},\dots,Y_{t},Y_{\mathsf{BExt}},\mathsf{BExt}(Z,Y_{\mathsf{BExt}})}}\right)$ $\approx_{\varepsilon_{\mathsf{block}}}\mathopen{}\mathclose{{\left(Y_{0},\dots,Y_{t},Y_{\mathsf{BExt}},\mathsf{BExt}(Z^{\prime\prime},Y_{\mathsf{BExt}})}}\right)$
    $\approx_{(t+1)\varepsilon_{0}}\mathopen{}\mathclose{{\left(Y_{0},\dots,Y_{t},Y_{\mathsf{BExt}},U_{m_{0}}}}\right).$